OpenStudy (anonymous):

what is the equation of the line containing the given pair of points (1/4,1/2) and (3/4,2)

OpenStudy (anonymous):

slope : (2-1/2) / (3/4 - 1/4) = (1.5) / ( 0.5) = 3 line equation : y = y1 + m(x-x1) y = 2+ 3(x-(3/4)) y = 2 + 3x - 9/4 y = -0.25 + 3x